OpenStudy (anonymous):

Find the distance between the points below. If necessary, round your answer to the nearest hundreth. (10,11) and (-2,2)?

hero (hero):

 Use pythagorean theorem

OpenStudy (anonymous):

(10,11) and (-2,2) to get the distance, you subtract x1 from x2, then y1 from y2 like this= 10 - (-2)=12 11-2=9 so the answer is (12,9)